OpenStudy (anonymous):

tell wether the equation is a solution to the linear equation (0,2);3x-2y=4,-2x-y=-2

OpenStudy (anonymous):

tell wether the equation is a solution to the linear equation (0,2);3x-2y=4,-2x-y=-2